A driver in a car traveling at a speed of 21.8 m/s sees a cat 101m away on the road. How long will it take for the car to accelerate uniformly to a stop in 99m?
1
Expert's answer
2011-10-13T13:01:23-0400
t-time, a -ecceleration
(1) at=21.8 m/s => t=21.8 /a
(2) at2/2=99m
237*62/a=99
a=2.4 m/s2
t=21.8/2.4=9.083 s